The larger of two numbers is one more than twice the smaller. The sum of the numbers is twenty less than three times the larger. Find the numbers.
Answer by mananth(16946)

let one number be x
the other number be y
.
y=2x+1
y-2x=1
.
x+y=3y-20
x-2y=-20
.
2(y-2x)+x-2y=2-20
2y-4x+x-2y=-18
-3x=-18
x= 6
y=2x+1
y=12+1
y=13

.